OpenStudy (anonymous):

solve for x x\[x^2 + \sqrt(3x) +3=0\]

OpenStudy (anonymous):

x\[x^2 + \sqrt(3x) +3=0\]

OpenStudy (anonymous):

help please???

OpenStudy (anonymous):

try raising everything to ^2 at first

OpenStudy (anonymous):

right then i get x^4 +3x+9=0

OpenStudy (anonymous):

this is where im stuck

OpenStudy (anonymous):

assume x2=y

OpenStudy (anonymous):

and replace

OpenStudy (anonymous):

y^2+3x+9=0

OpenStudy (anonymous):

actually do you know complex numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cause this is only complex. ok. from the beginning then. rewrite sqrt3x somehow else

OpenStudy (anonymous):

3x^.5

OpenStudy (anonymous):

nah. I guess sqrt3*sqrtx is comfier

OpenStudy (anonymous):

then, u get x^2+sqrt3*sqrtx=-3 which gives you 2 complex roots, cause a ^2 can't be negative in R